# Kusatsu, Shiga 525-8577, Japan

### v1

@@ 4.2 SCSMs with boundaries
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{ij}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}+\mu^{(1)}_{ia}(\bm{X}){\bm{Q}}^{i}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}\right],$ (117)
where $\mu^{(2)}$ is a pullback of a two-forms, $\mu^{(2)}\in\Omega^{2}(M)$, $\mu^{(1)}$ is a one-forms, $\mu^{(1)}$ taking a value in the pullback of $A$, and $\mu^{(0)}$ is a zero-forms, taking a value in pullback of $\wedge^{2}A$.
though we can consider more general terms including $p_{i},w_{a}$. The choice (117) is interesting since this choice has given interesting geometric structures including Dirac structures for ordinary CSMs [14]. and generalization of boundary conditions in gauged Poisson sigma models [6]. In the case of the CSM, boundary geometry is described by Dirac structures such as the twisted Poisson structure [36, 37]. For the case of gauged Poisson sigma models, boundary terms are described by momentum maps or momentum sections on Poisson manifolds [6].
@@
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
@@
$=2\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}\mathrm{d}{\bm{A}}^{a}+{\bm{Z}}_{i}{\bm{Q}}^{i}+\omega^{a}_{bi}{\bm{A}}^{b}{\bm{W}}_{a}{\bm{Q}}^{i}\right$.
$\left.+\rho^{i}_{a}(\bm{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{3!}H_{ijk}{\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}+\frac{1}{2}C_{ab}^{c}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right$.
@@
$\delta\bm{X}^{i}$ $={\{{{S},{\bm{X}^{i}}}\}}=\mathrm{d}\bm{X}^{i}-{\bm{Q}}^{i}-\rho^{i}_{a}{\bm{A}}^{a},$ (119)
@@
$=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-\frac{1}{3!}\partial_{[k}\mu^{(2)}_{ij]}{\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}-\frac{1}{2}(\rho^{k}_{a}\partial_{k}\mu^{(2)}_{ij}+\nabla_{[i}\mu_{j]a}^{(1)}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}\right$.
$\quad\left.-\left\{\left(\rho^{j}_{a}\partial_{j}\mu_{ib}^{(1)}-\frac{1}{2}C_{ab}^{c}\mu_{ic}^{(1)}\right)+\frac{1}{2}\nabla_{i}\mu^{(0)}_{ab}\right\}{\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}-\frac{1}{3!}\left(\rho^{i}_{[a}\partial_{i}\mu^{(0)}_{|bc]}+C_{[ab|}^{d}\mu^{(0)}_{d|c]}\right){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right],$ (122)
@@
There is a geometric interpretation of Eqs. (127)–(130). $\mu^{(0)}$, $\mu^{(1)}$ and $\mu^{(2)}$ satisfying (127)–(130) are a generalization of homotopy momentum sections on a (pre-)$2$-plectic manifold. In fact under Eq. (110) $\mu^{(0)}$, $\mu^{(1)}$ and $\mu^{(2)}$ satisfies the definition of homotopy momentum sections if $H^{(2)}=\iota_{\rho}H$, $H^{(1)}=\iota_{\rho}\iota_{\rho}H$, and $H^{(0)}=\iota_{\rho}\iota_{\rho}\iota_{\rho}H$ [29]. Moreover $\mu^{(0)}$, $\mu^{(1)}$ and $\mu^{(2)}$ reduces a homotopy moment map on a (pre-)multisympletic manifold with a Lie group action [28] if a Lie algebroid $A$ is an action Lie algebroid. If a Lie algebroid is $A=M\times\mathfrak{g}$, the $A$-differential $^{A}\mathrm{d}$ reduces to the Chevalley-Eilenberg differential of the Lie algebra $d_{CE}$ plus ${\mathcal{L}}_{\rho}$, and under some natural assumptions, Eqs. (127)–(130) reduces to the definition of a homotopy moment map [29]. Eqs. (127)–(130) are regarded as a generalization to the Courant algebroid setting.
@@ Acknowledgments
This work was supported by JSPS Grants-in-Aid for Scientific Research Number 22K03323.
@@ A.1 Definitions and examples
For a comprehensive review of Lie algebroids and their properties, see, for example, [39].
@@ A.2 Connections on Lie algebroids
as the dual connection. For a general discussion on connections on Lie algebroids, see [40, 41, 42].
@@
For two connections, $\nabla$ and $^{A}\nabla$, we introduce tensors: the curvature, $R\in\Omega^{2}(M,A\otimes A^{*})$, the $A$-torsion, $T\in\Gamma(A\otimes\wedge^{2}A^{*})$, and the basic curvature [43]. ${{}^{A}S}\in\Omega^{1}(M,\wedge^{2}A^{*}\otimes A)$. These are defined as follows:
@@ B.1 Definitions and examples
[4] A Courant algebroid is a vector bundle $E$ over $M$, which has a nondegenerate symmetric bilinear form $\langle-,\,$-$\rangle$, a bilinear operation $_{D}$} on $\Gamma(E)$ called the Dorfman bracket, and a bundle map called an anchor map, ${\rho_{(E)}}=\rho:E\longrightarrow TM$, satisfying the following properties:
@@ B.2 Connections on Courant algebroids
For a general discussion on connections on Courant algebroids, see [44, 45, 46]. For relations of curvatures, torsions, etc. of Courant algebroids with sigma models, refer to [47].
@@
for $e\in\Gamma(E)$, $s\in\Gamma(E^{\prime})$ and $f\in C^{\infty}(M)$ $^{E}\nabla_{fe}s=f{}^{E}\nabla_{e}s$ does not necessarily hold for an $E$-connection on a Courant algebroid. A proper definition of $E$-connection called a Dorfman connection appears in [45]. An improved definition of an $E$-connection which satisfying $\mathbb{R}$}-linearity is given by a Lie $2$-algebroid description of a Courant algebroid [48]. .
@@
For two connections, $\nabla$ and $^{E}\nabla$, we introduce important tensors: the curvature, $R\in\Omega^{2}(M,E\otimes E^{*})$, the $E$-torsion, or the generalized torsion, or the Gualtieri torsion, $T\in\Gamma(\wedge^{3}E^{*})$ [44], and the basic curvature ${{}^{E}S}\in\Omega^{1}(M,\wedge^{3}E^{*})$ [49]. These are defined as follows:
@@ C.1 Q-manifolds
There are some recent reviews of graded manifolds [3, 13, 15, 14] and references therein. We concentrate on nonnegatively graded manifolds called an N-manifold.
@@
Let $A$ be a vector bundle over $M$. AA is a Lie algebroid if and only if $A[1]$ is a Q-manifold with a homological vector field [50]. For a vector bundle $A$ over a smooth manifold $M$, $A[1]$ is a grade manifold such that degree of fiber coordinates is shifted by one. The concrete local expression is as follows. Let $(x^{i},a^{a})$ be local coordinates on $A[1]$ of degree $(0,1)$. Here $i$ is the index on the base manifold $M$ and $a$ is the index of the fiber. A vector field of degree +1+1 is locally given by
@@ C.2 QP-manifolds
[20] A QP-manifold of degree two has one to one correspondence to a Courant algebroid $E$.
@@ C.3 AKSZ sigma models
We briefly review sigma models called AKSZ sigma models [5]. For more details, refer to articles, for instance, [51, 3, 14, 52]
@@
Concrete construction is as follows. We assume a super differential $\mathrm{d}$ on $T[1]\Sigma$ induced from the de Rham differential on $\Sigma$ and a Berezin integration on $T[1]\Sigma$. Let $\bm{X}:\Sigma\rightarrow M$ for $\Sigma$ and the base manifold $M$ of $\mathcal{M}$}, Let $\vartheta$ be a Liouville $1$-form of ${\omega}_{grad}$ such that $\omega=-\delta\vartheta$.
@@
${\omega_{AKSZ}}$ $=\int_{T[1]\Sigma}\bm{X}^{*}{\omega}_{grad},$ (211)
@@
$S$ $=\int_{T[1]\Sigma}(\iota_{\mathrm{d}}\bm{X}^{*}\vartheta+\bm{X}^{*}\Theta),$ (212)
where $\bm{X}:T[1]\Sigma\rightarrow M$ is a map from $T[1]\Sigma$ from the degree zero base manifold $M$, and $\mathrm{d}$ is the super derivative on $T[1]\Sigma$. Local coordinates on $\mathcal{M}$} is mapped to corresponding superfields. We can prove that $S$ is a homological function, i.e., satisfies ${\{{{S},{S}}\}}=0$, which is the classical master equation in BV formalism. A brief interpretation of $S$ is that it generates a homological vector field induced from two differentials on $T[1]\Sigma$ and $\mathcal{M}$}, $\mathrm{d}+Q={\{{{S},{-}}\}}$.
@@ C.4 Manifestly covariant expressions
$x^{\prime i}=x^{i},$ (213)
$z^{\prime}_{i}=\frac{\partial x^{j}}{\partial x^{\prime i}}z_{j}+\frac{1}{2}k_{AB}M^{B}_{C}\partial_{i}M_{D}^{C}\eta^{A}\eta^{D},$ (214)
@@
Therefore a graded manifold $T^{*}[2]E[1]$ is not a cotangent bundle in the ordinary sense. The graded symplectic form is invariant under local coordinate transformations (213)–(215),
@@ D.1 Introduction of fluxes
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ija}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\bm{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (240)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\bm{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho_{a}^{i}{\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H_{ijk}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}+\frac{1}{2}T_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ija}(\bm{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\bm{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\bm{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (241)

### v2

@@ 4.2 SCSMs with boundaries
$S_{b}$ $=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[\frac{1}{2}\mu^{(2)}_{ij}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}+\mu^{(1)}_{ia}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{A}}^{a}+\frac{1}{2}\mu^{(0)}_{ab}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}\right],$ (117)
where $\mu^{(2)}$ is a pullback of a two-forms, $\mu^{(2)}\in\Omega^{2}(M)$, $\mu^{(1)}$ is a one-forms, $\mu^{(1)}$ taking a value in the pullback of $A^{*}$, and $\mu^{(0)}$ is a zero-forms, taking a value in pullback of $\wedge^{2}A^{*}$.
Though we can consider more general terms including $p_{i},w_{a}$, The choice (117) is interesting since this choice has given important geometric structures including Dirac structures for ordinary CSMs Ikeda:2012pv , and generalization of boundary conditions in gauged Poisson sigma models Ikeda:2023pdr . In the case of the CSM, boundary geometry is described by Dirac structures such as the twisted Poisson structure Klimcik:2001vg ; Park2000au . For the case of gauged Poisson sigma models, boundary terms are described by momentum maps or momentum sections on Poisson manifolds Ikeda:2023pdr .
@@
${\{{{S},{S}}\}}$ $=2\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\ \mathrm{d}\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}D{\bm{A}}^{a}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
@@
$=2\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-{\bm{Z}}_{i}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+{\bm{W}}_{a}\mathrm{d}{\bm{A}}^{a}+{\bm{Z}}_{i}{\bm{Q}}^{i}+\omega^{a}_{bi}{\bm{A}}^{b}{\bm{W}}_{a}{\bm{Q}}^{i}\right$.
$\left.+\rho^{i}_{a}(\boldsymbol{X}){\bm{Z}}_{i}{\bm{A}}^{a}+\frac{1}{3!}H_{ijk}{\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}+\frac{1}{2}C_{ab}^{c}{\bm{A}}^{a}{\bm{A}}^{b}{\bm{W}}_{c}\right$.
@@
$\delta\boldsymbol{X}^{i}$ $={\{{{S},{\boldsymbol{X}^{i}}}\}}=\mathrm{d}\boldsymbol{X}^{i}-{\bm{Q}}^{i}-\rho^{i}_{a}{\bm{A}}^{a},$ (119)
@@
$=\int_{T[1]\partial\Sigma}\mathrm{d}^{2}\sigma\mathrm{d}^{2}\theta\left[-\frac{1}{3!}\partial_{k}\mu^{(2)}_{ij}{\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}-\frac{1}{2}(\rho^{k}_{a}\partial_{k}\mu^{(2)}_{ij}+\nabla_{i}\mu_{ja}^{(1)}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}\right$.
$\quad\left.-\left\{\left(\rho^{j}_{a}\partial_{j}\mu_{ib}^{(1)}-\frac{1}{2}C_{ab}^{c}\mu_{ic}^{(1)}\right)+\frac{1}{2}\nabla_{i}\mu^{(0)}_{ab}\right\}{\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}-\frac{1}{3!}\left(\rho^{i}_{a}\partial_{i}\mu^{(0)}_{bc}+C_{ab}^{d}\mu^{(0)}_{dc}\right){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right],$ (122)
@@
There is a geometric interpretation of Eqs. (127)–(130). $\mu^{(0)}$, $\mu^{(1)}$ and $\mu^{(2)}$ satisfying (127)–(130) are a generalization of homotopy momentum sections on a (pre-)$2$-plectic manifold. In fact under Eq. (110) $\mu^{(0)}$, $\mu^{(1)}$ and $\mu^{(2)}$ satisfies the definition of homotopy momentum sections if $H^{(2)}=\iota_{\rho}H$, $H^{(1)}=\iota_{\rho}\iota_{\rho}H$, and $H^{(0)}=\iota_{\rho}\iota_{\rho}\iota_{\rho}H$ Hirota:2021isx . Moreover $\mu^{(0)}$, $\mu^{(1)}$ and $\mu^{(2)}$ reduces a homotopy moment map on a (pre-)multisympletic manifold with a Lie group action Callies:2013jbu if a Lie algebroid $A$ is an action Lie algebroid. If a Lie algebroid is $A=M\times\mathfrak{g}$, the $A$-differential $^{A}\mathrm{d}$ reduces to the Chevalley-Eilenberg differential of the Lie algebra $d_{CE}$ plus ${\mathcal{L}}_{\rho}$, and under some natural assumptions, Eqs. (127)–(130) reduces to the definition of a homotopy moment map Hirota:2021isx . Eqs. (127)–(130) are regarded as a generalization to the Courant algebroid setting.
@@ Acknowledgments
This work was supported by JSPS Grants-in-Aid for Scientific Research Number 22K03323 and by the research project, ’Higher structures in geometry and mathematical physics’ in the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center in Kyoto University.
@@ A.1 Definitions and examples
For a comprehensive review of Lie algebroids and their properties, see, for example, Mackenzie .
@@ A.2 Connections on Lie algebroids
as the dual connection. For a general discussion on connections on Lie algebroids, see DufourZung ; AbadCrainic ; CrainicFernandes .
@@
For two connections, $\nabla$ and $^{A}\nabla$, we introduce tensors: the curvature, $R\in\Omega^{2}(M,A\otimes A^{*})$, the $A$-torsion, $T\in\Gamma(A\otimes\wedge^{2}A^{*})$, and the basic curvature Blaom . ${{}^{A}S}\in\Omega^{1}(M,\wedge^{2}A^{*}\otimes A)$. These are defined as follows:
@@ B.1 Definitions and examples
LWX A Courant algebroid is a vector bundle $E$ over $M$, which has a nondegenerate symmetric bilinear form $\langle-,\,$-$\rangle$, a bilinear operation $_{D}$} on $\Gamma(E)$ called the Dorfman bracket, and a bundle map called an anchor map, ${\rho_{(E)}}=\rho:E\longrightarrow TM$, satisfying the following properties:
@@ B.2 Connections on Courant algebroids
For a general discussion on connections on Courant algebroids, see Gualtieri:2007bq ; Jotz ; Cueca:2019fej . For relations of curvatures, torsions, etc. of Courant algebroids with sigma models, refer to Chatzistavrakidis:2023lwo .
@@
[…]
@@
For two connections, $\nabla$ and $^{E}\nabla$, we introduce important tensors: the curvature, $R\in\Omega^{2}(M,E\otimes E^{*})$, the $E$-torsion, or the generalized torsion, or the Gualtieri torsion, $T\in\Gamma(\wedge^{3}E^{*})$ Gualtieri:2007bq , and the basic curvature ${{}^{E}S}\in\Omega^{1}(M,\wedge^{3}E^{*})$ Chatzistavrakidis:2023otk . These are defined as follows:
@@ C.1 Q-manifolds
There are some recent reviews of graded manifolds Roytenberg:2006qz ; Cattaneo:2010re ; Cueca:2025aej ; Ikeda:2012pv and references therein. We concentrate on nonnegatively graded manifolds called an N-manifold.
@@
Let $A$ be a vector bundle over $M$. AA is a Lie algebroid if and only if $A[1]$ is a Q-manifold with a homological vector field Vaintrob . For a vector bundle $A$ over a smooth manifold $M$, $A[1]$ is a grade manifold such that degree of fiber coordinates is shifted by one. The concrete local expression is as follows. Let $(x^{i},a^{a})$ be local coordinates on $A[1]$ of degree $(0,1)$. Here $i$ is the index on the base manifold $M$ and $a$ is the index of the fiber. A vector field of degree +1+1 is locally given by
@@ C.2 QP-manifolds
Roy01 A QP-manifold of degree two has one to one correspondence to a Courant algebroid $E$.
@@ C.3 AKSZ sigma models
We briefly review sigma models called AKSZ sigma models Alexandrov:1995kv . For more details, refer to articles, for instance, Cattaneo:2001ys ; Roytenberg:2006qz ; Ikeda:2012pv ; Ikeda:2025
@@
Concrete construction is as follows. We assume a super differential $\mathrm{d}$ on $T[1]\Sigma$ induced from the de Rham differential on $\Sigma$ and a Berezin integration on $T[1]\Sigma$. Let $\boldsymbol{X}:\Sigma\rightarrow M$ for $\Sigma$ and the base manifold $M$ of $\mathcal{M}$}, Let $\vartheta$ be a Liouville $1$-form of ${\omega}_{grad}$ such that $\omega=-\delta\vartheta$.
@@
${\omega_{AKSZ}}$ $=\int_{T[1]\Sigma}\boldsymbol{X}^{*}{\omega}_{grad},$ (211)
@@
$S$ $=\int_{T[1]\Sigma}(\iota_{\mathrm{d}}\boldsymbol{X}^{*}\vartheta+\boldsymbol{X}^{*}\Theta),$ (212)
where $\boldsymbol{X}:T[1]\Sigma\rightarrow M$ is a map from $T[1]\Sigma$ from the degree zero base manifold $M$, and $\mathrm{d}$ is the super derivative on $T[1]\Sigma$. Local coordinates on $\mathcal{M}$} is mapped to corresponding superfields. We can prove that $S$ is a homological function, i.e., satisfies ${\{{{S},{S}}\}}=0$, which is the classical master equation in BV formalism. A brief interpretation of $S$ is that it generates a homological vector field induced from two differentials on $T[1]\Sigma$ and $\mathcal{M}$}, $\mathrm{d}+Q={\{{{S},{-}}\}}$.
@@ C.4 Manifestly covariant expressions
$x^{\prime i}=x^{\prime i}(x),$ (213)
$z^{\prime}_{i}=\frac{\partial x^{j}}{\partial x^{\prime i}}z_{j}+\frac{1}{2}k_{AB}M^{B}_{C}\partial^{\prime}_{i}M_{D}^{C}\eta^{A}\eta^{D},$ (214)
@@
where $x^{\prime i}(x)$ is an arbitrary smooth functions of $x$. Therefore a graded manifold $T^{*}[2]E[1]$ is not a cotangent bundle in the ordinary sense. The graded symplectic form is invariant under local coordinate transformations (213)–(215),
@@ D.1 Introduction of fluxes
$S_{F}$ $=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[\frac{1}{2}H^{(2)}_{ija}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (240)
@@
$=\int_{T[1]\Sigma}\mathrm{d}^{3}\sigma\mathrm{d}^{3}\theta\left[-{\bm{Z}}_{i}^{\nabla}\mathrm{d}\boldsymbol{X}^{i}+{\bm{P}}_{i}\mathrm{d}{\bm{Q}}^{i}+\frac{1}{2}m_{ab}{\bm{A}}^{a}D{\bm{A}}^{b}+{\bm{Z}}_{i}^{\nabla}{\bm{Q}}^{i}+\rho_{a}^{i}{\bm{Z}}_{i}^{\nabla}{\bm{A}}^{a}\right$.
$\left.+\frac{1}{3!}H_{ijk}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{Q}}^{k}+\frac{1}{2}T_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right$.
$\left.+\frac{1}{2}H^{(2)}_{ija}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{Q}}^{j}{\bm{A}}^{a}+\frac{1}{2}H^{(1)}_{iab}(\boldsymbol{X}){\bm{Q}}^{i}{\bm{A}}^{a}{\bm{A}}^{b}+\frac{1}{3!}H^{(0)}_{abc}(\boldsymbol{X}){\bm{A}}^{a}{\bm{A}}^{b}{\bm{A}}^{c}\right].$ (241)